# Writer-Defined AI Personas for On-Demand Feedback Generation

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2023ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Text inputccs: Computing methodologies Natural language processing
@@ 2.2. Using Large Language Models for the Representation of People
Related work leveraged Natural Language Processing (NLP) to represent or “simulate” people (Park et al., 2023). A key line of work refers to “personas” as a foundation for generating text with language models from a certain perspective (e.g. (Song et al., 2019, 2021; Zhang et al., 2018; Zheng et al., 2020)), with HCI applications particularly for chatbots (e.g. (Haller and Rebedea, 2013; AMI, 2020; Schmitt and Buschek, 2021; Sun et al., 2023)). These personas are often short text descriptions. A related concept is the “system instruction” prompt, which tells an (instruction-tuned) model which perspective to take (e.g. “You are an assistant that speaks like Shakespeare” [^1] https://help.openai.com/en/articles/7042661-chatgpt-api-transition-guide).
@@
2.3. Writing Practice: Writing for an Audience
@@
2.4. Writing Research: Cognitive Processes of Writing
If writers should think about an inner reader, when do they think about the text? The different cognitive processes in writing are non-linear, as described by Flower and Hayes (Hayes and Flower, 1980; Flower and Hayes, 1981), involving planning, translating, and reviewing. Writers consider the target audience on several levels of planning, from concrete content goals and plans (e.g. what to say/do to an audience), to middle-range goals (e.g. “appealing to an audience”), to top-level goals (e.g. “interest the reader”). Writers work on ideas and plans towards these goals through a pattern of exploration and consolidation.
Our fundamental UI design accounts for this: Writers can easily switch from writing in a page view to a sidebar for capturing their understanding of the audience by creating and editing personas.
@@ 2.5. Designing Writing Tools: Expanding the Perspective Beyond the Writer
Writing tools have a long history (Macdonald et al., 1982; Kirschenbaum, 2016). NLP capabilities are increasingly integrated into them, both in industry (e.g. “Copilot” in MS Office (Spataro, 2023)), as well as in the HCI research community and its prototypes (e.g. see (Chang et al., 2023; Gero et al., 2022; Huang et al., 2022)).
@@ 3.1. Concept Development
3.1.1. Design Rationale
In summary, the related work motivated us to let writers define personas themselves, such that they can express their target readers’ characteristics, as well as obtain different perspectives. Writers can easily edit reader personas, to refine them, and also to react to potentially unwanted aspects in the generated feedback. They can change or define multiple personas, to support gaining insights through contrast, both into their text, as well as into the AI’s capability to manifest diversity in its generations. Finally, writers can request feedback on any piece of text (through text selection) to facilitate specificity of feedback.
@@ 3.2. Frontend
The frontend of our system is implemented as a web application using React.
@@ 3.2.1. Main layout: Editor and sidebar
The UI is split into two main areas – a text editor (Figure 1A) and a sidebar (Figure 1B). The sidebar’s navigation panel at the top consists of a “Feedback History” tab for requesting feedback, one tab for each created persona (to edit it), and a “+” button for adding a new persona.
@@ 3.2.2. Persona tabs: Create and edit personas
Each persona tab leads to a view with form fields (Figure 1B). Writers can give their persona a name that also shows up in relevant other places (tab header, feedback by that persona). The rest of the form is structured in rows of pairs of text fields (for attributes and descriptions, e.g. “job: physician”). Rows can be added and removed by clicking on respective buttons.
@@ 3.2.3. Feedback tab: View and request feedback from the personas
The “Feedback History” tab (Figure 1C) holds a list of interactive cards that display the generated feedback. New feedback is displayed at the top. Cards start in a collapsed state that cuts the text off after a few sentences. A “See more” button allows users to expand the card to view the full feedback. A “Show context” button at the bottom of the card can be clicked to show the part of the text that the feedback was based on.
@@ 3.2.4. Interaction
Figure 3. The overall workflow: (1) The writer defines a persona by adding rows of attributes and descriptions in a persona tab. (2) In the text editor, the writer selects a text part and then clicks on a persona button in the sidebar to generate feedback from that persona for that text.
@@
Figure 3. The overall workflow: (1) The writer defines a persona by adding rows of attributes and descriptions in a persona tab. (2) In the text editor, the writer selects a text part and then clicks on a persona button in the sidebar to generate feedback from that persona for that text.
@@
To request feedback from a created persona, writers first select the part of their text they want to receive feedback on. They then click on a persona’s button in the “Feedback History” tab. This adds a new comment card in the feedback history, displaying the feedback.
@@
If writers are not satisfied by a response, or identified extensions or new directions for the persona, they can switch to the persona’s tab at any time, to add, delete or edit fields.
@@ 3.3.2. Feedback Generation
We leverage OpenAI’s Chat Completion using GPT-3.5 to generate the persona feedback. See Section A.1 for our prompt template. We provided six input-output examples with this prompt template, as few-shot examples for the model. This was based on our observations in early experiments to improve the model’s adherence to the desired output. The feedback in these examples connected persona attributes to specific suggestions, to reveal how the effect of a persona’s attributes affects the generated feedback (e.g. “As a history professor, I value that this text…”). These six examples were created from our own interactions with the OpenAI model, based on interactive sessions in which we explored the model’s suitability at the start of the project (i.e. the few-shot examples are “cherry-picked” and manually refined input-outputs).
@@ 4.2. Participants
In total, 16 people participated in our evaluation: five in the first study (2 male, 3 female; referenced as $P_{1}$ $P_{5}$), using prototype V1; and eleven in the second one (7 male, 4 female; $P_{6}$ $P_{16}$), using V2. Their age ranged from 22-33 years. All were proficient English speakers; four were native speakers.
@@ 4.3.2. Writing Task (35 minutes)
We asked participants to share their screens with recording. We explained the task of working on their text. During the recruitment, we had asked participants to think of writing tasks that they could bring to the study. Eleven people brought an actual writing project, 5 had come up with one for the study. A researcher observed the task and took notes. They were muted for most of the time, occasionally reminding participants to share their thoughts or asking a question to better understand their thinking.
@@ 4.3.3. Interview (10 minutes)
We conducted semi-structured interviews after the writing task to further learn about people’s experiences. We asked these main questions: What was your thinking when creating the personas? What did you think about the generated feedback? How did you use the feedback? What did you like about this idea? What could be improved/extended?
@@ 4.3.4. Questionnaire (5 minutes)
Participants provided their demographics and in study 2 completed a final questionnaire asking about their experience with the overall concept as well as specific aspects and UI features. It included Likert items (see Figure 7) on Experience using the system and AI interest. We took the latter from the “MeMo:KI – Opinion Monitor AI” (mem, 2023).
@@
This figure shows a timeline. The timeline is divided into four parts: Study intro (10 min), Writing task (35 min), Interview (10 min), Questionnaire (5 min).
@@
4.4. Coding of Think-Aloud & Interviews
The transcribed interviews and observation notes were analyzed using a German qualitative content analysis method (Mayring, 2021) that is equivalent to thematic analysis (Braun and Clarke, 2006, 2019). Using axial coding principles (Corbin and Strauss, 2014), the first author and second author individually reviewed the texts several times, moving back and forth through the material. They identified overarching categories/themes and subcategories of codes through an iterative process of clustering, splitting, and merging codes. The final themes were then discussed with the other authors in videoconferencing sessions. We repeated this process until a consensus was reached.
@@ 5.1.1. Interaction (Logging Data)
We logged interactions in study 2: On average, participants created 2.36 personas (SD 0.48, range 2 to 3). They requested a mean of 6.27 feedback (SD 2.83, range 3 to 11), and revisited created personas 7.46 times on average (SD 5.70). Their final texts in the editor had a mean length of 438 words (SD 284.14).
@@ 5.2. How do writers perceive the idea of AI personas? (RQ1)
Related, $P_{12}$ contrasted the concept with AI tools that make text suggestions: “It’s much more valuable than just rewriting the sentence, because for that I could just use, ChatGPT, for example.”.
@@ 5.3.2. Interacting with Personas
Participants iteratively refined personas: They created a persona, requested feedback, and read that to decide whether it is the kind of feedback they wanted. If not, they added, removed or changed attributes. However, it was not always obvious how to achieve this. $P_{10}$ wondered: “What should I change? Is that level of granularity relevant?”. This workflow also motivated three participants to suggest conversations as a way to refine the persona, such as $P_{12}$: “Can I iterate the feedback? I’m missing a chat here to refine. […] It would be really cool to ask personas questions like ”would you use such a tool?””
@@ 5.3.3. Strategies for Writing with AI Personas
Figure 5 shows how participants split their time between working in the editor and the sidebar (based on logged UI events): Some switched only a few times, focusing on either writing or engaging with the personas for longer periods. In contrast, others switched frequently. Across both these groups, most participants started in the text editor (e.g. copy-pasting in the writing project they brought to our study) and then switched to creating personas during the first third of the study.
@@
Shows a vertical stack of horizontal bars, one for each participant. Each bar is colored in stripes that represent the participant’s focus switches over time, between the text editor and the sidebar. Additional lines indicate the moments at which a new persona was added. The emerging patterns reveal that some participants switched between the two UI parts more frequently than others.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642406isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Text inputccs: Computing methodologies Natural language processing
@@ 2.2. Using Large Language Models for the Representation of People
Related work leveraged Natural Language Processing (NLP) to represent or “simulate” people (Park et al., 2023). A key line of work refers to “personas” as a foundation for generating text with language models from a certain perspective (e.g. (Song et al., 2019, 2021; Zhang et al., 2018; Zheng et al., 2020)), with HCI applications particularly for chatbots (e.g. (Haller and Rebedea, 2013; AMI, 2020; Schmitt and Buschek, 2021; Sun et al., 2023a)). These personas are often short text descriptions. A related concept is the “system instruction” prompt, which tells an (instruction-tuned) model which perspective to take (e.g. “You are an assistant that speaks like Shakespeare” [^1] https://help.openai.com/en/articles/7042661-chatgpt-api-transition-guide).
@@
2.3. Writing Research & Practice: Writing for an Audience
@@
If writers should think about an inner reader, when do they think about the text? The different cognitive processes in writing are non-linear, as described by Flower and Hayes (Hayes and Flower, 1980; Flower and Hayes, 1981), involving planning, translating, and reviewing. Writers consider the target audience on several levels of planning, from concrete content goals and plans (e.g. what to say/do to an audience), to middle-range goals (e.g. “appealing to an audience”), to top-level goals (e.g. “interest the reader”). Writers work on ideas and plans towards these goals through a pattern of exploration and consolidation. Our fundamental UI design accounts for this: Writers can easily switch from writing in a page view to a sidebar for capturing their understanding of the audience by creating and editing personas.
@@ 2.4. Designing Writing Tools: Expanding the Perspective Beyond the Writer
Writing tools have a long history (Macdonald et al., 1982; Kirschenbaum, 2016). NLP capabilities are increasingly integrated into them, both in industry (e.g. “Copilot” in MS Office (Spataro, 2023)), as well as in the HCI research community and its prototypes (e.g. see (Chang et al., 2023; Gero et al., 2022; Huang et al., 2022; Lee et al., 2024)).
@@ 3.1.1. Design Iteration
3.1.2. Design Rationale and Goals
We extracted the following higher-level design rationale and goals from the literature and insights from our first user test with V1: 1) As the key idea, supporting on-demand feedback allows writers to receive feedback from a reader perspective while writing, through AI personas. 2) Enabling writers to define personas themselves allows them to express their target readers’ characteristics, and to obtain different perspectives in this way. 3) Related, supporting easy iteration of personas enables writers to refine them and also to react to potentially unwanted aspects in the generated feedback. 4) Supporting the definition and comparison of multiple personas facilitates gaining insights through contrast, both into the draft, as well as into the AI’s capability to manifest diversity in its generations. 5) Finally, supporting local feedback allows writers to request feedback on any piece of text (which we realized through text selection) to facilitate specificity of feedback.
@@ 3.2. Frontend
The frontend of our system is implemented as a web application using React to achieve the mentioned design goals.
@@ 3.2.1. Main layout: Editor and sidebar
The UI is split into two main areas – a text editor (Figure 1A) and a sidebar (Figure 1B). The sidebar’s navigation panel at the top consists of a “Feedback History” tab for requesting feedback, one tab for each created persona (to edit it), and a “+” button for adding a new persona. This side-by-side layout supports the mentioned design goals via easy access to both the draft and the personas.
@@ 3.2.2. Persona tabs: Create and edit personas
This tab view supports the two design goals of enabling writers to define personas themselves and to easily iterate on personas. Each persona tab leads to a view with form fields (Figure 1B). Writers can give their persona a name that also shows up in relevant other places (tab header, feedback by that persona). The rest of the form is structured in rows of pairs of text fields (for attributes and descriptions, e.g. “job: physician”). Rows can be added and removed by clicking on respective buttons.
@@ 3.2.3. Feedback tab: View and request feedback from the personas
The “Feedback History” tab (Figure 1C) realizes the main design goal of providing on-demand feedback, as well as the goal of allowing for comparisons of feedback from multiple personas. It holds a list of interactive cards that display the generated feedback. New feedback is displayed at the top. Cards start in a collapsed state that cuts off the text after a few sentences. A “See more” button allows users to expand the card to view the full feedback. A “Show context” button at the bottom of the card can be clicked to show the part of the text that the feedback was based on.
@@ 3.2.4. Interaction
Figure 3. The overall workflow: (1) The writer defines a persona by adding rows of attributes and descriptions in a persona tab. (2) In the text editor, the writer selects a part of the text and then clicks on a persona button in the sidebar to generate feedback from that persona for that piece of text.
@@
Figure 3. The overall workflow: (1) The writer defines a persona by adding rows of attributes and descriptions in a persona tab. (2) In the text editor, the writer selects a part of the text and then clicks on a persona button in the sidebar to generate feedback from that persona for that piece of text.
@@
To request feedback from a created persona, writers first select the part of their text they want to receive feedback on (in line with the design goal of supporting local feedback). They then click on a persona’s button in the “Feedback History” tab. This adds a new comment card in the feedback history, displaying the feedback.
@@
If writers are not satisfied by a response, or identified extensions or new directions for the persona, they can switch to the persona’s tab at any time, to add, delete or edit fields; this supports easy iteration.
@@ 3.3.2. Feedback Generation
We leverage OpenAI’s Chat Completion using GPT-3.5 to generate the persona feedback. See Section A.1 for our prompt template. We provided six input-output examples with this prompt template, as few-shot examples for the model. This was based on our observations in early experiments to improve the model’s adherence to the desired output. The feedback in these examples connected persona attributes to specific suggestions, to reveal how a persona’s attributes affect the generated feedback (e.g. “As a history professor, I value that this text…”). These six examples were created from our own interactions with the OpenAI model, based on interactive sessions in which we explored the model’s suitability at the start of the project (i.e. the few-shot examples are “cherry-picked” and manually refined input-outputs).
@@ 4.2. Participants
In total, 16 people participated in our evaluation (Table 1): five in the first study (2 male, 3 female; referenced as $P_{1}$ $P_{5}$), using prototype V1; and eleven in the second one (7 male, 4 female; $P_{6}$ $P_{16}$), using V2. Their age ranged from 22-33 years. All were proficient English speakers; four were native speakers.
@@
Table 1. Overview of the participants. *Writing proficiency, **English proficiency (both on Likert scales, 1/low to 5/high).
P Writing Project Regular Audiences Writing* English** Age Gender Occupation
P1
Bachelor thesis, Email to Professor, Fiction Story
Creative Writing, Video Game Writing, Scientific Writing
4
5
25
Female
Student
P2
Blog post (AI in education)
P2 themselves, professors
4
5
26
Female
Student
P3
Research Paper (HCI)
Scientific community
4
5
32
Male
Researcher
P4
Research Grant Proposal (Computer Science)
Researchers, general public, students
4
4
33
Male
Associate Professor
P5
Blog post (Cooking)
P5 themselves, professors
3
3
32
Female
Student
P6
Web article (Computer Architecture)
Academic audience
5
5
23
Male
Student
P7
Business Report (ERP system in HR)
Managers, Co-workers, professors
2
4
22
Male
IT-Consultant
P8
Research Paper (HCI)
Scientific reviewers, Management Teams, C-Level of coporates/start-ups
4
4
29
Female
PhD equivalent
P9
Research Paper (HCI)
Human-Computer Interaction and Human-Robot Interaction audiences
4
5
33
Male
PhD equivalent
P10
Research Paper (HCI)
Domain experts, researchers, colleagues
5
5
32
Male
PhD equivalent
P11
Research Paper (Process Analytics)
Computer Science scientists
3
5
27
Male
PhD equivalent
P12
Research Paper (HCI)
Scientific conferences, academia, work-related audiences
5
5
25
Male
PhD equivalent
P13
Research Paper (NLP)
Academic audiences in the fields of NLP, Machine Learning, etc.
4
5
23
Male
PhD equivalent
P14
Masters Thesis Abstract (AI in medicine)
None
2
3
27
Female
Patient management in hospital
P15
Research Paper Summary (Canine psychology)
Uni (exams, course work, presentations)
3
5
31
Female
Student
P16
Research Essay (English language variety)
Professors at university
3
5
26
Female
Student
This table gives an overview of the participants including columns that describe the writing project, regular audiences, writing proficiency, English proficiency, age, gender, and occupation for each participant.
Table 1. Overview of the participants. *Writing proficiency, **English proficiency (both on Likert scales, 1/low to 5/high).
@@ 4.3.2. Writing Task (35 minutes)
We asked participants to share their screens with recording. We explained the task of working on their text. During the recruitment, we had asked participants to think of writing tasks that they could bring to the study. Eleven people brought an actual writing project, 5 had come up with one for the study. A researcher observed the task and took notes. They were muted for most of the time, occasionally reminding participants to share their thoughts or asking a question to better understand people’s thinking.
@@ 4.3.3. Interview (10 minutes)
We conducted semi-structured interviews after the writing task to further learn about people’s experiences. We asked these main questions: “What was your thinking when creating the personas?”, “What did you think about the generated feedback?”, “How did you use the feedback?”, “What did you like about this idea?”, “What could be improved/extended?”
@@ 4.3.4. Questionnaire (5 minutes)
Participants provided their demographics and in study 2 completed a final questionnaire asking about their experience with the overall concept as well as specific aspects and UI features. It included Likert items on Experience using the system and AI interest. We took the latter from the “MeMo:KI – Opinion Monitor AI” (mem, 2023). These Likert items were constructed as statements reflecting the participant’s viewpoint, such as “The feedback influenced my writing”, “The feedback made it easier for me to write my text”, and “I read articles about Artificial Intelligence with great curiosity”. See Figure 8 for all items.
@@
This figure shows a timeline. The timeline is divided into four parts: Study intro (10 minutes), Writing task (35 minutes), Interview (10 minutes), Questionnaire (5 minutes).
@@
4.4. Coding of Think-Aloud, Interviews, and the Generated Feedback
The transcribed interviews and observation notes were analyzed using a German qualitative content analysis method (Mayring, 2021) that is equivalent to thematic analysis (Braun and Clarke, 2006, 2019). Using axial coding principles (Corbin and Strauss, 2014), the first author and second author individually reviewed the texts several times, moving back and forth through the material. They identified overarching categories/themes and subcategories of codes through an iterative process of clustering, splitting, and merging codes. The final themes were then discussed with the other authors in videoconferencing sessions. We repeated this process until a consensus was reached. In the same way, we also coded the generated feedback texts.
@@ 5.1.1. Interaction (Logging Data)
We logged interactions in study 2: On average, participants created 2.36 personas (SD 0.48, range 2 to 3). They requested a mean of 6.27 feedbacks (SD 2.83, range 3 to 11), and revisited created personas 7.46 times on average (SD 5.70). Their final texts in the editor had a mean length of 438 words (SD 284.14). Figure 4(b) shows an overview of the attribute-description pairs that participants entered to define their personas.
Figure 5. Overview of the words that participants entered in the attribute-description pairs (see UI in Figure 2), lightly scaled by how often these contributed to generated feedback (also annotated as numbers if ¿1). For example, if a participant defined “writing style: formal, scientific” for a persona and requested feedback four times from this persona, this would contribute counts of four to “writing” and “style” (in 4(a)), as well as to “formal” and “scientific” (in 4(b)). We removed stopwords (e.g. “the”). While shown as single words here for a better overview, participants also entered descriptions as phrases, not only single words (e.g. “Provide real-world analogies to ideas in the paper”). Log files including verbatim descriptions are available in the project repository (see link in Section 7).
(a) Attribute words used in personas.
(b) Description words used in personas.
Figure with two separate word clouds: The left one shows the words that participants entered as attributes of personas (top words: “content”, “writing”, “style”, “role”, “task”). The right one shows the same for the words that participants entered as descriptions (top words: “paper”, “research”, “formal”, “sentences”, “computer”).
Figure 5. Overview of the words that participants entered in the attribute-description pairs (see UI in Figure 2), lightly scaled by how often these contributed to generated feedback (also annotated as numbers if ¿1). For example, if a participant defined “writing style: formal, scientific” for a persona and requested feedback four times from this persona, this would contribute counts of four to “writing” and “style” (in 4(a)), as well as to “formal” and “scientific” (in 4(b)). We removed stopwords (e.g. “the”). While shown as single words here for a better overview, participants also entered descriptions as phrases, not only single words (e.g. “Provide real-world analogies to ideas in the paper”). Log files including verbatim descriptions are available in the project repository (see link in Section 7).
@@ 5.1.2. Perception of Interaction (Likert Data)
5.1.3. Overview of the Generated Feedback
By manually coding and analyzing the 68 generated feedback texts (Section 4.4), we identified a recurring structure with the following building blocks: In the opening sentences, the persona described its attributes, typically like this: “As a [e.g., reviewer, professor, journalist, …], with [certain attributes], my task is […].” This was almost always followed by a positive comment or summary of the user’s selected text snippet, and often concluded with a suggestion on how to improve the text.
In the feedback’s main part, the AI persona suggested adding more examples (47 occurrences), asked for topic-related content (26), sought clarifications (34) and more details (13), or gave specific hints, such as missing data, highlights, need for citations, or conclusions. In addition, the persona proposed improvements to the writing style (9), including suggestions for shorter sentences, easier language for accessibility, or different terminology. This advice was typically first described on a higher level, followed by a concrete suggestion (27), most often with a variation of “For example, instead of [text snippet], the author could write [proposed text].”
Finally, almost all feedback texts (66) concluded with a summary, often starting with “Overall, the text snippet…” or less frequently with phrases like “In summary…” or “In conclusion…”.
Generally, we observed the AI personas to provide polite and uplifting feedback. Many times (63 occurrences), the feedback emphasized positive aspects of the text. Even when one participant ($P_{12}$) created a “mean” persona, although less encouraging, the feedback remained polite and included self-critical statements like “[…] the mean attitude in this context may affect the credibility of the feedback.”
While the feedback’s structure consistently followed the described scheme, its content was tailored to each participant’s text snippet and the persona’s preferences and background, rather than offering generic suggestions (e.g. see Figure 7).
@@ 5.2. How do writers perceive the idea of AI personas? (RQ1)
Five people drew explicit comparisons to other AI tools. For example, $P_{12}$ contrasted the concept with AI tools that give text suggestions: “It’s much more valuable than just rewriting the sentence, because for that I could just use, ChatGPT, for example.”. Similarly, $P_{15}$ said: “[the persona interaction] feels more like a person, like you kind of have this picture in your head of who you’re talking to. It’s a bit different with like generic AI models like ChatGPT […].” When asked about what he liked about the concept, $P_{13}$ highlighted that “this layout and like being able to add personas […] it’s like a nice layer over whatever language model you’re using underneath”. These comments point at the perceived usefulness of the persona concept beyond LLM-based support in general.
@@ 5.3.2. Interacting with Personas
Participants iteratively refined personas: They created a persona, requested feedback, and read that to decide whether it is the kind of feedback they wanted. If not, they added, removed or changed attributes. However, it was not always obvious how to achieve this. $P_{10}$ wondered: “What should I change? Is that level of granularity relevant?”. This workflow also motivated three participants to suggest conversations as a way to refine the persona, such as $P_{12}$: “Can I iterate the feedback? I’m missing a chat here to refine. […] It would be really cool to ask personas questions like would you use such a tool?
@@ 5.3.3. Strategies for Writing with AI Personas
Figure 6 shows how participants split their time between working in the editor and the sidebar (based on logged UI events): Some switched only a few times, focusing on either writing or engaging with the personas for longer periods. In contrast, others switched frequently. Across both these groups, most participants started in the text editor (e.g. copy-pasting in the writing project they brought to our study) and then switched to creating personas during the first third of the study. After engaging with their personas for the first time, participants asked for feedback approximately every 3 minutes and 37 seconds. The most engaged participant sought feedback every two minutes, while the least engaged waited 4 minutes and 45 seconds. For participants who switched more frequently, two patterns emerge: Some (4) seek feedback mainly from various personas on the same text snippet, while others (2) consistently use a single persona across different text parts. We found no relationship between frequency of use and the perceived helpfulness of personas.
@@
Shows a vertical stack of horizontal bars, one for each participant. Each bar is colored in stripes that represent the participant’s focus switches over time, between the text editor and the sidebar. Additional lines indicate the moments at which a new persona was added. The emerging patterns reveal that some participants switched between the two user interface parts more frequently than others.